# Intelligent System for Automated Molecular Patent Infringement Assessment

### v1

@@ 2.3 Evaluation of MarkushParser
To demonstrate the effectiveness of MarkushParser in parsing Markush images into strings, we conduct experiments on MolPatent-240. We maintain the condition that the query molecules are always input in SMILES format and compared three different Markush input methods: original patent PDFs, Markush images, and Markush strings parsed by MarkushParser. We utilize three LLMs that support vision inputs – GPT-4o, Claude, and Gemini – to meticulously compare their intermediate outputs and the process of chain-of-thought reasoning under the different input modalities. The experimental results are shown in Table 1 and Fig.5a. The figure illustrates that the LLMs struggle to accurately understand the structure of Markush under the image and PDF understanding schemes, often erroneously fabricating Markush structures and non-existent substituent groups for analysis.
In Fig.5b, we present several results of the MarkushParser’s parsing of Markush images. The first row in the figure displays the original input Markush images, while the second row shows the extended SMILES strings generated by the MarkushParser and the reconstructed Markush structures based on these parsing results. The MarkushParser effectively captures the structural details of the Markush structures, and the reconstructed outputs exhibit high consistency with the original inputs.
@@ 4.1.1 Construction of MolPatent-240
Due to the lack of publicly available molecular patent infringement datasets, we manually construct MolPatent-240 to facilitate evaluation. This dataset comprises 240 patent–molecule pairs, each labeled as either “protected” or “not protected” by human experts. To construct this dataset, we collect 70 molecular protection patents from public online sources, primarily including patents from WIPO (the World Intellectual Property Organization, prefixed with “WO”) and USPTO (the United States Patent and Trademark Office, prefixed with “US”). For positive examples, we select existing embodiments from the patents or manually designed novel molecules based on the patent claims. For negative examples, we modify the structures of positive examples at the atomic level, introducing subtle differences in the molecular skeleton or substituents that deviate from the patent claims. To ensure data accuracy, the dataset underwent multiple rounds of manual evaluation. After construction, three experts specializing in drug discovery verify the chemical correctness of all samples. This includes validating the legality of the constructed molecules and the correctness of their protection labels. Based on their feedback, individual dataset contributors revise the benchmark to correct errors and remove a subset of samples that are too challenging to rectify. The final dataset consists of 240 samples, including 104 positive examples (protected samples) and 136 negative examples (non-protected samples).
@@ 4.1.2 Construction of Substituent Groups Extraction Data
Training MarkushMatcher requires large-scale data, which is impractical to annotate manually with human experts. Furthermore, the task of matching Markush structures with molecules and extracting substituent groups is exceptionally challenging, as rule-based algorithms and LLMs fail to produce satisfactory results. To address this, we propose a reverse data construction method that samples Markush structures and substituent groups from large-scale databases, assembles them into complete molecules using chemical tools, and records the substituent matching relationships during the assembly process. The entire pipeline is depicted in Fig.4.
MolParser-7M 333https://huggingface.co/datasets/AI4Industry/MolParser-7M, a dataset containing over 7 million molecular and structural images, serves as the foundation for this process. Some samples in MolParser-7M include Markush structures. We heuristically design filtering rules to exclude unsuitable samples, such as those without Markush structures, those containing multiple molecules, or molecules that are excessively small. This filtering process results in a dataset of 847 thousand Markush structure records. We manually curate and organize a substituent group database containing 4.7k substituents derived from molecular structures. The distribution of the substituent groups in the training set is shown in Fig.3e. For each Markush sample in the dataset, we randomly sample substituent groups from this database according to the number of substituent attachment points on the scaffold. Using RDKit, we combine the Markush scaffold with the sampled substituents based on predefined rules to generate complete molecules. The substituent matching relationships are used as supervisory signals during the model’s training procedure.
Through this reverse construction method, we generate a dataset containing 847 thousand structural matching data. We randomly sample 1,000 examples from this dataset for model evaluation, with the results shown in Figure 3. The remaining samples are used to train MarkushMatcher.
@@ 4.1.3 Construction of Markush OCSR Data
To facilitate the parsing of Markush diagrams in the patent, we collect 515 real-world Markush images, and manually label their corresponding string expressions. During labeling, we employ the extended SMILES notation defined in [47] to represent Markush structures as text. This dataset includes hand-drawn molecular images, Markush structures extracted from molecular protection patents, and other noisy samples.
@@ 4.2 LLM-based Agents
Skeleton Extractor. This agent is responsible for extracting the protected molecular skeletons from the patent. It supports reading patent content from both online sources and local PDF files with Patent Web Search and deployed parsing tool, respectively. It identifies key protected Markush structures and the associated claim requirements, specifically the substituent group values outlined in the patent.
Substitutes Matcher. This agent assesses the structural similarity between the query molecule and the extracted molecular skeleton to determine skeleton matches. It is equipped with a substructure matching algorithm and a substituent group extraction model. Both results from the matching algorithm and the extraction model are provided as references to this agent. Using vision-LLM (GPT-4o with image input), this agent verifies substituent group mappings, corrects chemical errors made during the substituent group extraction, and performs necessary rearrangements to the mappings based on the requirements.
Requirements Examiner. This agent compares the query molecule with the requirements defined in the patent, conducting the analysis at both the skeleton level and the substituent group level. At the skeleton level, the agent assesses the query molecule against the Markush structural skeleton and relevant requirements outlined in the patent (e.g., the overall number of heavy atoms or whether the molecule is explicitly discussed in the patent). For the substituent group level comparison, the agent evaluates the values of the substituent groups on the molecule against the definitions provided in the patent. After completing these comparisons, the agent reports the results to the Planner.
@@ 4.2.1 Neural Network Tool Models
MarkushParser. MarkushParser converts molecular images into extended SMILES representations with OCSR (Optical Chemical Structure Recognition) and is capable of translating common molecules, substituent groups (incomplete molecules with connection points), and Markush structures.
@@
We initialize MarkushParser with MolParser [47], which is pre-trained on a large annotated dataset consisting of 7.3 million samples, encompassing both real-world and synthetic image-SMILES pairs of molecules and Markush. We manually collect and label 515 Markush images and their associated extended SMILES strings from real-world patent or drug discovery scenarios. We train the model using supervised fine-tuning (SFT) with our 515 Markush image-string pairs for 20 epochs with a learning rate of .
@@ 4.2.2 Heuristic Tools
Patent Web Search. This function retrieves segmented molecular images and plain text content from the Google Patents database 444https://patents.google.com/. It is particularly useful for handling input queries when only the patent ID is provided. Notably, our framework remains fully functional without internet-based support, as it can alternatively rely on local PDF parsing and image processing through the OCSR scheme. This flexibility ensures that users can still perform comprehensive analyses even when web-based resources are unavailable.

### v2

@@ 2.3 Evaluation of MarkushParser
To demonstrate the effectiveness of MarkushParser in parsing Markush images into strings, we conduct experiments on MolPatent-240. We maintain the condition that the query molecules are always input in SMILES format and compare three different Markush input methods: original patent PDFs, Markush images, and Markush strings parsed by MarkushParser. We utilize three LLMs that support vision inputs – GPT-4o, Claude, and Gemini – to meticulously compare their intermediate outputs and the process of chain-of-thought reasoning under the different input modalities. The experimental results are shown in Table 1 and Figure5a. The figure illustrates that the LLMs struggle to accurately understand the structure of Markush under the image and PDF understanding schemes, often erroneously fabricating Markush structures and non-existent substituent groups for analysis.
In Figure5b, we present several results of the MarkushParser’s parsing of Markush images. The first row in the figure displays the original input Markush images, while the second row shows the extended SMILES strings generated by the MarkushParser and the reconstructed Markush structures based on these parsing results. The MarkushParser effectively captures the structural details of the Markush structures, and the reconstructed outputs exhibit high consistency with the original inputs.
@@ 4.1.1 Construction of MolPatent-240
Due to the lack of publicly available molecular patent infringement datasets, we manually construct MolPatent-240 to facilitate evaluation. This dataset comprises 240 patent–molecule pairs, each labeled as either “protected” or “not protected” by human experts.
To construct this dataset, we collect 70 molecular protection patents from public online sources. For positive examples, we select existing embodiments from the patents or manually designed novel molecules based on the patent claims. For negative examples, we modify the structures of positive examples at the atomic level, introducing subtle differences in the molecular skeleton or substituents that deviate from the patent claims.
After construction, three human experts specializing in drug discovery verify the chemical correctness of all samples. This includes validating the legality of the constructed molecules and the correctness of their protection labels. Based on their feedback, individual dataset contributors revise the benchmark to correct errors and remove a subset of samples that are too challenging to rectify.
@@ 4.1.2 Construction of Substituent Groups Extraction Data
Training MarkushMatcher requires large-scale data, which is impractical to annotate manually with human experts. Furthermore, the task of matching Markush structures with molecules and extracting substituent groups is exceptionally challenging, as rule-based algorithms and LLMs fail to produce satisfactory results. To address this, we propose a reverse data construction method that samples Markush structures and substituent groups from large-scale databases, assembles them into complete molecules using chemical tools, and records the substituent matching relationships during the assembly process. The entire pipeline is depicted in Figure4.
MolParser-7M [41], a dataset containing over 7 million molecular and structural images, serves as the foundation for this process. Some samples in MolParser-7M include Markush structures. We heuristically design filtering rules to exclude unsuitable samples, such as those without Markush structures, those containing multiple molecules, or molecules that are excessively small. This filtering process results in a dataset of 847 thousand Markush structure records. We manually curate and organize a substituent group database containing 4.7 thousand substituents derived from molecular structures. The distribution of the substituent groups in the training set is shown in Figure3e. For each Markush sample in the dataset, we randomly sample substituent groups from this database according to the number of substituent attachment points on the scaffold. Using RDKit, we combine the Markush scaffold with the sampled substituents based on predefined rules to generate complete molecules. The substituent matching relationships are used as supervisory signals during the model’s training procedure.
Through this reverse construction method, we generate a dataset containing 847 thousand pieces of structural matching data. We randomly sample 1,000 examples from this dataset for model evaluation, with the results shown in Figure 3. The remaining samples are used to train MarkushMatcher.
@@ 4.1.3 Construction of Markush OCSR Data
To facilitate the parsing of Markush diagrams in the patent, we collect 515 real-world Markush images and manually label their corresponding string expressions. During labeling, we employ the extended SMILES notation defined in [41] to represent Markush structures as text. This dataset includes hand-drawn molecular images, Markush structures extracted from molecular protection patents, and other noisy samples.
@@ 4.2 LLM-based Agents
Sketch Extractor. This agent is responsible for extracting the protected molecular skeletons from the patent. It supports reading patent content from both online sources and local PDF files with Patent Web Search and deployed parsing tool, respectively. It identifies key protected Markush structures and the associated claim requirements, specifically the substituent group values outlined in the patent.
Substituents Matcher. This agent assesses the structural similarity between the query molecule and the extracted molecular skeleton to determine skeleton matches. It is equipped with a substructure matching algorithm and a substituent group extraction model. Both results from the matching algorithm and the extraction model are provided as references to this agent. Using vision-LLM (GPT-4o with image input), this agent verifies substituent group mappings, corrects chemical errors made during the substituent group extraction, and performs necessary rearrangements to the mappings based on the requirements.
Requirements Examinator. This agent compares the query molecule with the requirements defined in the patent, conducting the analysis at both the skeleton level and the substituent group level. At the skeleton level, the agent assesses the query molecule against the Markush structural skeleton and relevant requirements outlined in the patent (e.g., the overall number of heavy atoms or whether the molecule is explicitly discussed in the patent). For the substituent group level comparison, the agent evaluates the values of the substituent groups on the molecule against the definitions provided in the patent. After completing these comparisons, the agent reports the results to the Planner.
@@ 4.2.1 Neural Network Tool Models
MarkushParser. MarkushParser converts molecular images into extended SMILES representations with OCSR (Optical Chemical Structure Recognition) and can translate common molecules, substituent groups (incomplete molecules with connection points), and Markush structures.
@@
We manually collect and label 515 high-quality Markush images and their associated extended SMILES strings from real-world patent or drug discovery scenarios. The model is first initialized with MolParser [41], then fine-tuned with our 515 Markush image-string pairs for 20 epochs with a learning rate of .
@@ 4.2.2 Heuristic Tools
Patent Web Search. This function retrieves segmented molecular images and plain text content from the Google Patents database 333https://patents.google.com/. It is particularly useful for handling input queries when only the patent ID is provided. Notably, our framework remains fully functional without internet-based support, as it can alternatively rely on local PDF parsing and image processing through the OCSR scheme. This flexibility ensures that users can still perform comprehensive analyses even when web-based resources are unavailable.
@@
Data availability
During the training data construction of MarkushMatcher, we use the open-source Markush databank at https://huggingface.co/datasets/AI4Industry/MolParser-7M. Our patent infringement assessment benchmark MolPatent-240 is released at https://github.com/syr-cn/patentfinder_code_private/tree/master/patent_finder/data.
Code availability
The code of PatentFinder is available at https://github.com/syr-cn/patentfinder_code_private, including the code of the multi-agent inference framework, implementations of the heuristic tools, and neural network tools. The 847 thousand constructed data for training MarkushMatcher and the trained model parameters are available at https://osf.io/ftq9g. The 515 human-labeled data used for fine-tune MarkushParser are available at https://osf.io/jsqa3.
Corresponding authors
Correspondence authors: Xiang Wang and Hengxing Cai.
Competing interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix A Implementation Details
Within the PatentFinder framework, LLMs are assigned distinct roles and collaborate systematically to address the task of patent infringement analysis. During the process, input prompts are carefully tailored to optimize the models’ adaptation to their respective roles. This section outlines the role-specific prompts designed to guide the cognitive processes of the models.
Initially, each agent is equipped with the foundational information necessary to comprehend the definition and structure of extended SMILES strings.
Subsequently, each intelligent agent is equipped with a carefully crafted, human-written prompt tailored to its specific role within the system.
Appendix B Statistics of MolPatent-240
MolPatent-240 consists of 240 molecule-patent pairs, including 104 positive examples (protected samples) and 136 negative examples (non-protected samples). Positive samples are either directly selected from patent embodiments or manually constructed to meet specific patent requirements. Negative samples are created by violating patent substituent requirements, altering the core Markush skeleton, or selecting embodiments from unrelated patents. Table 2 summarizes the number of samples for each label type and the corresponding construction approach.
| Label | #Samples | Construction Method |
| Positive | 67 | Selected from embodiments in the patents. |
| Positive | 37 | Manually constructed based on the patent requirements. |
| Negative | 73 | Manually constructed by violating the patent substituent requirements. |
| Negative | 20 | Manually constructed by violating the core Markush skeleton. |
| Negative | 43 | Selected from embodiments in unrelated patents. |
| Total | 240 | - |
The dataset comprises 70 unique patents, with 9 sourced from WIPO (World Intellectual Property Organization, prefixed with “WO”) and 61 from USPTO (United States Patent and Trademark Office, prefixed with “US”). The publication years of these patents, as illustrated in Figure 6, span the past 20 years, with over 90% being recent patents published within the last decade.
To enhance the diversity of MolPatent-240, we deliberately curated patents containing a wide array of Markush structures. For visualization purposes, the core Markush skeletons from these 70 patents were extracted, and their Morgan fingerprints computed. The t-SNE dimensionality reduction of these fingerprints is presented in Figure 7, where each point represents the core Markush structure of a patent. The size of each point corresponds to the number of heavy atoms in the Markush structure. This visualization highlights substantial variation in the features of the claimed Markush skeletons, with a relatively uniform distribution in terms of both size and structural characteristics. The diversity of the MolPatent-240 benchmark ensures a robust and comprehensive evaluation of different algorithms.
Appendix C Fine-Grained Case Study
In Fig. 2, we present a case study illustrating the analytical process of PatentFinder and an example of its generated infringement report. For brevity, the displayed infringement report has been manually edited to fit within the layout constraints. In this section, we provide the model’s raw output, which includes more detailed information, such as intermediate outputs from each intelligent agent and the final summarized patent infringement report.
The key information from the patent, as extracted by the Sketch Extractor, is as follows:
The substructure match between the Markush claim and the query molecule, as identified by the Substituents Matcher, is as follows:
The detailed examination of the patent’s requirements, as conducted by the Requirements Examinator, is as follows:
The analysis of the patent PDF block indices that substantiate the above reasoning, as provided by the Fact Checker, is as follows:
The final infringement report, summarized by the Infringement Reporter, is as follows:
Appendix D Reproducibility
When applying multi-agent systems to scientific discovery tasks, a crucial aspect is evaluating the reproducibility of their outputs to ensure stability in real-world applications. To assess the reproducibility of PatentFinder, we randomly select a sample from the MolPatent-240 dataset and conducted five independent patent infringement analyses. As illustrated in Figure 8, all substituents comply with the definitions provided in the corresponding patent and are therefore protected under its claims. The primary challenges in this case include: (1) The Markush structure predominantly contains undefined substituents, posing significant challenges for structural recognition and match prediction by the tool model. (2) The patent specifies distinct definitions for the and sides, resulting in special cases mentioned in Figure 3f and 3g within the Markush structure. These require the language model to integrate patent content with substituent-matching reasoning.
For the experiments, we employed OpenAI-o1 as the backbone language model of PatentFinder, with a temperature of . Figure 8 illustrates the five patent infringement reports generated during the analysis. The results show that in 4 out of 5 cases, PatentFinder successfully matched the molecular structure with the patent’s claim requirements, arriving at the correct conclusion. However, in the fourth report, an error occurred when PatentFinder misjudged the substituent as non-compliant with the requirement of being an “optionally substituted 4- to 6-membered heterocyclic ring having 1-2 heteroatoms”. This issue arises from limitations in the language model’s understanding of molecular structures. Consequently, the performance of PatentFinder can be further enhanced through advancements in the backbone language model and by adapting scientific language models. Additionally, a simple majority voting strategy can improve the stability of PatentFinder. By generating multiple infringement reports and deciding the final prediction based on majority votes, the system can achieve greater reliability. For instance, in this case, PatentFinder could correctly predict the molecule as being protected by the patent in 4 out of 5 analyses, yielding the correct outcome.
Appendix E Bad Case Analysis
Experiments on the MolPatent-240 dataset demonstrate that while PatentFinder surpasses baseline models in overall accuracy for patent infringement prediction, it is not yet capable of perfectly resolving all molecular infringement cases. To better understand the limitations of PatentFinder and guide future research, we analyze two failed predictions from MolPatent-240, as illustrated in Figure 9.
Sample a. The molecular structure in this example perfectly matches the Markush scaffold, and its substituents meet all patent requirements. Therefore, it is supposed to be classified as protected. During PatentFinder’s reasoning process, the MarkushMatcher module correctly identifies the substituent matches. However, the Substituents Matcher module erroneously modifies the match results during the correction step, ultimately misclassifying the molecule as not protected. This reveals a tendency of the language model to overconfidently override correct predictions provided by specialized submodules.
Sample b. This molecule contains a sulfur-containing heterocyclic ring (highlighted in red area) that does not match the corresponding structure in the Markush claim, and it should therefore be classified as not protected. However, the Requirements Examinator module incorrectly concluded that the query molecule’s scaffold matches the Markush structure. This misjudgment led to subsequent errors in substituent matching, resulting in a false positive prediction that the molecule is protected. This highlights the model’s limitations in analyzing complex chemical structures.
Analysis of these failure cases reveals that while PatentFinder can successfully conduct parts of its reasoning process and remains faithful to the patent’s original requirements, certain steps still lead to errors. Despite the incorrect conclusions, the intermediate reasoning steps can provide valuable insights for drug discovery researchers, showcasing the high interpretability and potential utility of the PatentFinder framework.
Appendix F Limitations
While PatentFinder demonstrates promising performance in patent infringement analysis, surpassing traditional algorithms and baseline methods relying solely on text-based language models, several challenges remain for its deployment in drug discovery workflows: (1) Instability Introduced by LLMs: Despite the multi-agent reasoning system and the incorporation of domain-specific tool models designed to mitigate hallucinations and enhance factual accuracy, LLMs can still produce responses with logical flaws or factual inaccuracies. Neither the multi-agent system nor the tool models can entirely eliminate the possibility of errors during the LLM’s reasoning process, posing a risk to the robustness of the framework. (2) Difficulty in Evaluating the Analytical Process: Patent infringement assessment involves complex reasoning, and evaluating each intermediate step of the multi-agent system is challenging. Although this paper introduces the MolPatent-240 dataset to assess model performance, the evaluation is limited to treating patent infringement determination as a binary classification task. Due to the inherent complexity of patent analysis, it is difficult to establish ground truth for each intermediate step and design appropriate evaluation metrics. This limitation makes further assessment of PatentFinder highly dependent on manual labor, constraining the speed and scale of evaluations. (3) Dependence on Upstream Patent Retrieval Systems: PatentFinder requires relevant patents to be gathered by an upstream patent retrieval system for comparison. Errors in this retrieval process – such as failing to provide patents pertinent to the target molecule – may result in incorrect conclusions, such as classifying a protected molecule as unprotected. Such errors could lead to significant legal risks.
Despite its advancements, PatentFinder cannot fully replace human effort in this domain. However, with ongoing progress in artificial intelligence, more powerful backbone language models and advanced molecular analysis tools are expected to enhance PatentFinder’s capabilities in patent analysis and molecular understanding. These advancements will further unlock the potential of its multi-agent architecture, improving both performance and reliability in patent infringement analysis and related applications.